OpenStudy (anonymous):

what length of soled cylinder 2 cm in diamatre must be taken to recast into a ahollow cylinder of external diametre 20cm,0.25cm thick and 15cm long?

OpenStudy (anonymous):

@kittiwitti1

OpenStudy (kittiwitti1):

Oh. Uh.... Can't help you with that. e_e @hartnn

OpenStudy (anonymous):

why cany u help me with this??

OpenStudy (anonymous):

cant*

OpenStudy (kittiwitti1):

I suck at geometry.

OpenStudy (anonymous):

ok?

OpenStudy (kittiwitti1):

I'm gonna afk for 20 minutes while I try to solve this. e_e;

OpenStudy (anonymous):

...

OpenStudy (kittiwitti1):

I told you I suck at geometry....

OpenStudy (anonymous):

thanks anyways

OpenStudy (kittiwitti1):

np e_e

OpenStudy (kittiwitti1):

When you say "0.25 cm thick" do you mean width?

OpenStudy (anonymous):

idk

OpenStudy (kittiwitti1):

....how am I supposed to give you the correct answer then

OpenStudy (anonymous):

i haven't understood the sum

OpenStudy (kittiwitti1):

Sum?

OpenStudy (kittiwitti1):

 does this sound accurate?

OpenStudy (perl):

so the solid cylinder is melted down and recast into a hollow cylinder

OpenStudy (perl):

and you want know how much solid cylinder (in length) do you need to melt down and recast into the hollow cylinder.

OpenStudy (kittiwitti1):

I think I got the answer... this is when I wish I had a draw tablet e_e;

OpenStudy (perl):

helps to draw the hollow cylinder

OpenStudy (kittiwitti1):

I did on my paper. xD

OpenStudy (anonymous):

?

OpenStudy (kittiwitti1):

You know what, listen to @perl... she's the expert xD (You are a she right e_e?)

OpenStudy (perl):

I am using formula Volume = Pi * r^2 * h

OpenStudy (perl):



OpenStudy (anonymous):

now?

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

100-64?

OpenStudy (anonymous):

24*15?

OpenStudy (anonymous):

thats no answer

OpenStudy (anonymous):

in my book its 74.06

OpenStudy (perl):

there was a typo

OpenStudy (anonymous):

?

OpenStudy (perl):

Volume of solid cylinder formula : V = Pi * radius^2 * height substitute: V1 = pi * (1)^2 * h Volume of hollow cylinder: formula : V = Pi * (outer_radius^2 - inner_radius^2 ) * height substitute V2 = pi * [ 10^2 - (10 - .25)^2 ] * 15 From the problem we must have V1 = V2 Pi * (1)^2 * h = pi * (10^2 - (10 - .25)^2 )*15 divide both sides by Pi 1* h = [ 10^2 - (10 - .25)^2 ] * 15 h = ( 10^2 - (10 - .25)^2) * 15 h = ( 10^2 - 9.75^2 ) * 15 h = ( 100 - `95.0625`) * 15 h = `4.9375` * 15 h = 74.0625